# thanks: Corresponding author

### v1

@@ VI Summary and conclusion
In the present study, we performed a comprehensive analysis of the proton electromagnetic FFs using EP measurements in kinematic regions where the BH process dominates the cross section. Considering the strong dependence of the BH amplitude on the Dirac and Pauli FFs, $F_{1}(t)$ and $F_{2}(t)$, we examined several fitting scenarios based on dipole and $P$-pole parametrizations and extracted them as well as the corresponding Sachs FFs, $G_{E}(t)$ and $G_{M}(t)$.
Our results indicate that the EP data in the range $0.11<|t|<0.45~\text{GeV}^{2}$ provide crucial constraints on the Dirac form factor $F_{1}(t)$, whereas the sensitivity to the Pauli form factor $F_{2}(t)$ remains limited due to the reduced impact of its contribution at small values of $|t|$.
@@
Overall, the present study indicates that EP process, when measured in regions dominated by the BH process, provide a powerful and complementary tool for constraining the electromagnetic structure of the proton. Extending EP measurements to even smaller values of $|t|$ allow for a more precise determination of the nucleon radii and may offer additional insight into the long-standing charge-radius puzzle. The analysis framework developed in this study provides a solid foundation for future combined studies including both EP and elastic electron-nucleon scattering data which enables a unified extraction of the nucleon electromagnetic FFs.

### v2

@@ VI Summary and conclusion
In the present study, based on CLAS measurements of exclusive photon leptoproduction, we performed a comprehensive analysis of the proton electromagnetic FFs in kinematic regions where the BH process dominates the cross section. Considering the strong dependence of the BH amplitude on the Dirac and Pauli FFs, $F_{1}(t)$ and $F_{2}(t)$, we examined several fitting scenarios based on dipole and $P$-pole parametrizations and extracted them as well as the corresponding Sachs FFs, $G_{E}(t)$ and $G_{M}(t)$.
Our results indicate that the EP data in the range $0.11<|t|<0.45~\text{GeV}^{2}$ can provide constraints on the Dirac form factor $F_{1}(t)$, whereas the sensitivity to the Pauli form factor $F_{2}(t)$ remains limited due to the reduced impact of its contribution at small values of $|t|$.
@@
Overall, the present study indicates that EP process, when measured in regions dominated by the BH process, provide a complementary tool for constraining the electromagnetic structure of the proton. Extending EP measurements to even smaller values of $|t|$ allow for a more precise determination of the nucleon radii and may offer additional insight into the long-standing charge-radius puzzle. The analysis framework developed in this study provides a solid foundation for future combined studies including both EP and elastic electron-nucleon scattering data which enables a unified extraction of the nucleon electromagnetic FFs.
In addition, the methodology developed in this work can be directly applied to other exclusive photon leptoproduction measurements, such as those from Hall A Defurne and others (2015, 2017), provided that sufficiently precise cross-section data are available in kinematic regions where the BH contribution is dominant. This offers promising opportunities for future studies with expanded kinematic coverage and improved precision.